# Stochastic Thermodynamics of Non-reciprocally Interacting Particles and Fields

### v2

@@ A Transition affinities
Importantly, $\varepsilon^{int},s^{b},s^{state}$ is a property of the lattice, thus a cumulative change in it due to multiple stochastic transitions only depends on the initial and final lattice state. This subsequently applies to $\varepsilon$ and $\varepsilon^{stoch}$. In comparison, $\Delta_{i}^{\vec{\mathcal{D}}\#}f^{nr}$ and $\Delta_{\gamma\gamma^{\prime}}^{\#}f^{nr}$ are defined locally on the lattice. The stochastic transition entropy is defined as:
@@
The eq. 11 quantifies the thermodynamic cost of each microscopic transition supported by the environment. We adhere to the convention ‘forces generate currents’ throughout this paper, where, forces refer to the transition affinities, for instance, $\Delta\varepsilon,\Delta s^{state},\Delta f^{nr},\vec{f}_{i}^{sp},f_{\gamma\gamma^{\prime}}^{ch}$. The transition affinities are categorized into boundary (conservative) terms $(\Delta\varepsilon,\Delta s^{state})$ and bulk (non-conservative) terms $(\Delta f^{nr},\vec{f}_{i}^{sp},f_{\gamma\gamma^{\prime}}^{ch})$, which correspond to the relaxation and dissipative currents respectively [20]. Considering a set of consecutive stochastic transitions in observation time $\tau$, the conservative and non-conservative total entropy production due to all transitions is $O(1)$ and $O(\tau)$ respectively. Importantly, the non-conservative forces fundamentally differ in their origin. $\Delta f^{nr}$ and $f_{\gamma\gamma^{\prime}}^{ch}$ (or $\vec{f}_{i}^{sp}$) lie in the state-space and transition-space respectively. Moreover, $f_{i}^{nr}$ depends on the local particle occupancy, unlike $f_{\gamma\gamma^{\prime}}^{ch}$ (or $\vec{f}_{i}^{sp}$) being constants. Thus, the non-conservative driving due to the non-reciprocal interactions depends on the initial and final lattice states. The non-conservative driving due to non-reciprocal interactions is more dynamic in comparison to fixed enzymatic or self-propulsion driving forces.
@@ B Microscopic EPR
The microscopic reactive and diffusive mean EPR are defined as $\langle\dot{\sigma}^{r}\rangle$ and $\langle\dot{\sigma}^{d}\rangle$ respectively [1, 20]:
@@
Here, $\sum_{\{\Delta_{\gamma\gamma^{\prime}}^{\#}\}}^{\{N\}}$ and $\sum_{\{\Delta_{i}^{\vec{\mathcal{D}}\#}\}}^{\{N\}}$ denotes sum over all lattice configurations for the reactive and diffusive transitions respectively. It satisfies the fundamental definition of EPR namely force times current, $\langle\dot{\sigma}^{r}\rangle=\sum_{\{\Delta_{\gamma\gamma^{\prime}}^{\#}\}}^{\{N\}}j_{\gamma\gamma^{\prime}}^{\#}(\{N\})\Delta_{\gamma\gamma^{\prime}}^{\#}\sigma$ and $\langle\dot{\sigma}^{d}\rangle=\sum_{\{\Delta_{i}^{\vec{\mathcal{D}}\#}\}}^{\{N\}}j_{i}^{{\vec{\mathcal{D}}\#}}(\{N\})\Delta_{i}^{\vec{\mathcal{D}}\#}\sigma$ due to the LDB eq. 5 and the definitions eqs. 9 and 11. $\langle\dot{\sigma}^{r}\rangle>0$ and $\langle\dot{\sigma}^{d}\rangle>0$ due to the inequality $(x-y)\ln{\left(x/y\right)}$ leads to the second law of thermodynamics. The total mean microscopic EPR reads $\langle\dot{\sigma}\rangle=\langle\dot{\sigma}^{r}\rangle+\langle\dot{\sigma}^{d}\rangle$.
@@ C Conservative and non-conservative decomposition of EPR
We reorganize $\langle\dot{\sigma}\rangle$ using eqs. 7, 8 and 5, obtaining the EPR contributions due to the reciprocal, non-reciprocal, Gibbs entropic and external non-conservative driving forces.
@@
We define the mean microscopic work rate $\langle\dot{w}\rangle=-\sum_{\{N\}}P_{\{N\}}\partial_{t}\varepsilon^{int}$ due to the explicit time-dependent driving of the control parameters $\{\lambda\}$ of $\varepsilon^{int}$ [19]. Hence, $d_{t}\langle\varepsilon\rangle=\sum_{\{N\}}\partial_{t}P_{\{N\}}\varepsilon^{int}+P_{\{N\}}\partial_{t}\varepsilon^{int}$ incorporates the stochastic work. The Gibbs entropy and Gibbs EPR reads [91]:
[…]
@@
The eq. 13 is the second law of thermodynamics decomposed on the basis of the origin of the forces acting along the transition. The first two terms are derived from the conservative forces. Because, their time-integrated EPR is given by change of a functional defined for the lattice, therefore it depends only on the initial and final state. In contrast, the remaining terms give EPR due to the non-conservative forces, namely non-reciprocal, self-propulsion and enzymatic chemical driving.
@@ D Orthogonal decomposition of the EPR
We define the Boltzmann probability distribution $P^{\varepsilon}_{\{N\}}(t)=e^{-\varepsilon+\psi_{\varepsilon}}$ with reference energy functional $\varepsilon$. We compute the free energy $\psi_{\varepsilon}=-\log{(\mathcal{Z}^{\varepsilon})}$ in terms of the partition function $\mathcal{Z}^{\varepsilon}=\sum_{\{N\}}e^{-\varepsilon}$. In the absence of the non-conservative forces, i.e. $f_{i}^{nr},\vec{f}_{i}^{sp},f_{\gamma\gamma^{\prime}}^{ch}=0$, the system satisfies the Boltzmann distribution. We decompose eq. 10 using $P^{\varepsilon}_{\{N\}}(t)$ to [94, 95, 96]:
@@
Comparing eq. 17 to eq. 10, analogous to eq. 9 we define the reference state entropy $s^{state}_{\varepsilon}$,
@@
Where, $-D_{\varepsilon}^{KL}(t)=\langle s^{state}_{\varepsilon}\rangle$ holds. Analogous to $d_{t}s^{gb}$, the rate of change of $D_{\varepsilon}^{KL}(t)$ is defined as:
@@
We introduce the orthogonal decomposition of the $\langle\dot{\sigma}^{ch}\rangle$ and $\langle\dot{\sigma}^{sp}\rangle$,
@@
Here, $j_{\gamma\gamma^{\prime}}^{ch}$ is the anti-symmetric part of $j_{\gamma\gamma^{\prime}}^{\#}(\{N\})$ under the adjoint transformation $f_{\gamma\gamma^{\prime}}^{ch}\to-f_{\gamma\gamma^{\prime}}^{ch}$. Similarly, $j_{i}^{sp}$ is the anti-symmetric part of $j_{i}^{\vec{\mathcal{D}}\#}(\{N\})$ under the adjoint transformation $\vec{f}_{i}^{sp}\to-\vec{f}_{i}^{sp}$. We have dropped the ##\## and $\{N\}$ index for the brevity. The exact expression for $j_{\gamma\gamma^{\prime}}^{ch}$ and $j_{i}^{sp}$ reads:
@@
In eq. 23, we have decomposed the EPR into its four linearly independent orthogonal contributions. First, $-d_{t}\psi_{\varepsilon}$ quantifies the rate of change of the free energy attributed to the external driving work needed to change the control parameters $\{\lambda\}$ of $\varepsilon$. Second, $-d_{t}D^{KL}_{\varepsilon}(t)$ quantifies the EPR due to the relaxation towards the Boltzmann distribution $P_{\{N\}}^{\varepsilon}$. Where, $-d_{t}\psi_{\varepsilon}$ and $-d_{t}D^{KL}_{\varepsilon}(t)$ are physically interpreted as the boundary terms in the $\{\lambda\}$ and $P_{\{N\}}$ space respectively. Third, $\langle\dot{\sigma}^{nr}\rangle$ quantifies the EPR due to the anti-symmetric forces (and vorticity currents generated by it) between the non-reciprocal particles. Fourth, $\langle\dot{\sigma}^{sp}\rangle+\langle\dot{\sigma}^{ch}\rangle$ quantifies the EPR due to the external non-conservative force ‘along the transition’.
One identifies the non-adiabatic EPR $\langle\dot{\sigma}_{\varepsilon}^{na}\rangle=-d_{t}D^{KL}_{\varepsilon}(t)$ and the remaining terms as the housekeeping EPR $\langle\dot{\sigma}_{\varepsilon}^{hs}\rangle$. The orthogonal decomposition is more fundamental because it does not require the existence of a steady state, as $\varepsilon$ is well-defined for any dynamic state. The lower bound on the total EP is obtained by using $\langle\dot{\sigma}_{\varepsilon}^{na}\rangle$ [101, 98], it reads $\langle\sigma_{\varepsilon}^{na}\rangle\geq D^{KL}_{\varepsilon}(0)-D^{KL}_{\varepsilon}(\tau)\geq D^{KL}(P(0)||P(\tau))$. Other notable consequences of the orthogonal decomposition have utilized different reference gauges [102, 103, 104, 105, 106, 107, 108, 109]. By choosing $P_{\{N\}}^{ss}$ as a reference distribution, the non-adiabatic housekeeping decomposition of $\langle\dot{\sigma}\rangle$ is detailed in appendix A.
@@ 2.4 Coarse-graining
We implement the Doi-Peliti coarse-graining (DPCG) procedure to obtain coarse-grained description [110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123]. The DPCG procedure incorporates the fluctuations (discreetness of the microscopic particle nature) due to the second quantization approach used. The technical details of the coarse-graining procedure and its physical implications are summarized in Ref. [86].
@@ 3 Fluctuating Macroscopic Description
The set of coarse-grained mesostates are denoted by $\{N_{i}^{\#}\}$. Where, $\{N_{i}^{\#}\}$ is defined over the discrete (lattice) space for the mesoscopic description. Physically, it corresponds to lumping the microstates (particles) with equivalent dynamic and thermodynamic properties to the extensive coarse-grained mesostate. It achieves the coarse-graining step from (a) to (b) shown in fig. 1. We define $\Omega=N_{tot}/\mathcal{V}$, where $N_{tot}$ and $\mathcal{V}$ are the total number of particles and the volume of the lattice (number of lattice sites), respectively. Thus, $\Omega$ quantifies the mean number of particles per lattice site. $N_{i}^{\#}$ is $O(\Omega)$ and its fluctuations due to the microscopic transitions are $O(1)$ [124, 86].
To proceed further with the mesoscopic to macroscopic coarse-graining step, (b) to (c) depicted in fig. 1, we introduce the mesoscopic to macroscopic scaling $N_{i}^{\#}=\Omega\rho_{i}(\vec{x})$ between the macrostate $\rho_{i}(x)$ and the particle number mesostate $N_{i}^{\#}$. Therefore, $\rho_{i}$ and its fluctuations due to microscopic transitions are $O(1)$ and $O(1/\Omega)$ respectively [124, 86]. By construction, $\Omega\geq 1$, as $\Omega<1$ contradicts the mesoscopic to macroscopic coarse-graining. The space field $\vec{x}$ tracks the lattice index, which we drop in our notations and its meaning is presumed. Importantly, we use the macrostate convention consistent with the large-deviation theory [124, 86]. Where, an intensive variable is defined and its fluctuations are suppressed with the large-deviation scaling parameter $\Omega$. Thus, the notion of an intensive variable is defined in the context of the fluctuations. The macrostate $\rho_{i}$ does not necessarily coincide with the convention of the density used in hydrodynamics unless $\Omega=\mathcal{V}$ is chosen. This is fulfilled in the hydrodynamic limit, where the average number of particles per lattice site scales with the system volume. Hence, $N_{i}^{\#}=\mathcal{V}\rho_{i}$ holds. Importantly, this flexible definition of the macrostate $\rho_{i}$ allows us to define and study the systems on intermediate observation scales between the mesoscopic and hydrodynamic descriptions. $\Omega=1$ corresponds to the mesoscopic description that utilizes the particle number as the coarse-grained state. The thermodynamic limit of $\Omega\to\infty$ suppresses macrostate fluctuations and recovers the deterministic limit corresponding to the reaction-diffusion systems, fig. 1(d).

### v3

@@ A Transition affinities
Importantly, $\varepsilon^{int},s^{b},s^{state}$ are properties of the lattice; therefore, a cumulative change in them due to multiple stochastic transitions depends only on the initial and final lattice states. This subsequently applies to $\varepsilon$ and $\varepsilon^{stoch}$. In comparison, $\Delta_{i}^{\vec{\mathcal{D}}\#}f^{nr}$ and $\Delta_{\gamma\gamma^{\prime}}^{\#}f^{nr}$ are defined locally on the lattice. The stochastic transition entropy is defined as:
@@
Equation 11 quantifies the thermodynamic cost of each microscopic transition supported by the environment. We adhere to the convention ‘forces generate currents’ throughout this paper, where forces refer to the transition affinities, for instance, $\Delta\varepsilon,\Delta s^{state},\Delta f^{nr},\vec{f}_{i}^{sp},f_{\gamma\gamma^{\prime}}^{ch}$. Transition affinities are categorized into boundary (conservative) terms $(\Delta\varepsilon,\Delta s^{state})$ and bulk (non-conservative) terms $(\Delta f^{nr},\vec{f}_{i}^{sp},f_{\gamma\gamma^{\prime}}^{ch})$, which correspond to relaxation and dissipative currents, respectively [20]. Considering a set of consecutive stochastic transitions over observation time $\tau$, the conservative and non-conservative total entropy production due to all transitions is $O(1)$ and $O(\tau)$, respectively. Importantly, the non-conservative forces fundamentally differ in their origin. $\Delta f^{nr}$ and $f_{\gamma\gamma^{\prime}}^{ch}$ (or $\vec{f}_{i}^{sp}$) lie in the state-space and the transition-space, respectively. Moreover, $f_{i}^{nr}$ depends on the local particle occupancy, unlike $f_{\gamma\gamma^{\prime}}^{ch}$ (or $\vec{f}_{i}^{sp}$), which are constants. Thus, the non-conservative driving due to the non-reciprocal interactions depends on the initial and final lattice states. The non-conservative driving due to non-reciprocal interactions is more dynamic in comparison to fixed enzymatic or self-propulsion driving forces.
@@ B Microscopic EPR
The microscopic reactive and diffusive mean EPR are defined as $\langle\dot{\sigma}^{r}\rangle$ and $\langle\dot{\sigma}^{d}\rangle$, respectively [1, 20]:
@@
Here, $\sum_{\{\Delta_{\gamma\gamma^{\prime}}^{\#}\}}^{\{N\}}$ and $\sum_{\{\Delta_{i}^{\vec{\mathcal{D}}\#}\}}^{\{N\}}$ denote the sum over all lattice configurations for the reactive and diffusive transitions, respectively. This satisfies the fundamental definition of mean EPR, namely, force times current: $\langle\dot{\sigma}^{r}\rangle=\sum_{\{\Delta_{\gamma\gamma^{\prime}}^{\#}\}}^{\{N\}}j_{\gamma\gamma^{\prime}}^{\#}(\{N\})\Delta_{\gamma\gamma^{\prime}}^{\#}\sigma$ and $\langle\dot{\sigma}^{d}\rangle=\sum_{\{\Delta_{i}^{\vec{\mathcal{D}}\#}\}}^{\{N\}}j_{i}^{{\vec{\mathcal{D}}\#}}(\{N\})\Delta_{i}^{\vec{\mathcal{D}}\#}\sigma$, due to the LDB condition eq. 5 and the definitions eqs. 9 and 11. $\langle\dot{\sigma}^{r}\rangle>0$ and $\langle\dot{\sigma}^{d}\rangle>0$ due to the inequality $(x-y)\ln{\left(x/y\right)}$, which leads to the second law of thermodynamics. The total mean microscopic EPR reads $\langle\dot{\sigma}\rangle=\langle\dot{\sigma}^{r}\rangle+\langle\dot{\sigma}^{d}\rangle$.
@@ C Conservative and non-conservative decomposition of EPR
We reorganize $\langle\dot{\sigma}\rangle$ using eqs. 7, 8 and 5, obtaining the EPR contributions due to the reciprocal, non-reciprocal, Gibbs entropic, and external non-conservative driving forces.
@@
Here, $d_{t}\langle\varepsilon\rangle$ and $d_{t}s^{gb}$ are attributed to the rate of change of $\varepsilon^{int}-s^{b}$ and $s^{state}$ from eq. 11. We define the mean microscopic work rate $\langle\dot{w}\rangle=-\sum_{\{N\}}P_{\{N\}}\partial_{t}\varepsilon^{int}$ due to the explicit time-dependent driving of the control parameters $\{\lambda\}$ of $\varepsilon^{int}$ [19]. Hence, $d_{t}\langle\varepsilon\rangle=\sum_{\{N\}}\partial_{t}P_{\{N\}}\varepsilon^{int}+P_{\{N\}}\partial_{t}\varepsilon^{int}$ incorporates the stochastic work. The Gibbs entropy and the Gibbs EPR read [96]:
$\begin{split}s^{gb}=-\sum_{\{N\}}P_{\{N\}}\ln{\left(P_{\{N\}}\right)},\hskip 14.22636ptd_{t}s^{gb}=-\sum_{\{N\}}d_{t}P_{\{N\}}\ln{\left(P_{\{N\}}\right)}.\end{split}$ (14)
@@
Equation 13 is the second law of thermodynamics decomposed on the basis of the origin of the forces acting along the transitions. The first two terms are derived from conservative forces. Because their time-integrated EPR is given by the change of a functional defined for the lattice, it depends only on the initial and final states. In contrast, the remaining terms give the EPR due to the non-conservative forces, namely non-reciprocal, self-propulsion, and enzymatic chemical driving.
@@ D Orthogonal decomposition of the EPR
We define the Boltzmann probability distribution $P^{\varepsilon}_{\{N\}}(t)=e^{-\varepsilon+\psi_{\varepsilon}}$ with reference energy functional $\varepsilon$. We compute the free energy $\psi_{\varepsilon}=-\log{(\mathcal{Z}^{\varepsilon})}$ in terms of the partition function $\mathcal{Z}^{\varepsilon}=\sum_{\{N\}}e^{-\varepsilon}$. In the absence of non-conservative forces, i.e., $f_{i}^{nr},\vec{f}_{i}^{sp},f_{\gamma\gamma^{\prime}}^{ch}=0$, the system satisfies the Boltzmann distribution. We decompose eq. 10 using $P^{\varepsilon}_{\{N\}}(t)$ to [99, 100, 101]:
@@
Comparing eq. 17 to eq. 10, analogously to eq. 9, we define the reference state entropy $s^{state}_{\varepsilon}$,
@@
Where $D_{\varepsilon}^{KL}(t)=\langle s^{state}_{\varepsilon}\rangle$ holds. Analogously to $d_{t}s^{gb}$, the rate of change of $D_{\varepsilon}^{KL}(t)$ is defined as:
@@
We introduce the orthogonal decomposition of $\langle\dot{\sigma}^{ch}\rangle$ and $\langle\dot{\sigma}^{sp}\rangle$,
@@
Here, $j_{\gamma\gamma^{\prime}}^{ch}$ is the anti-symmetric part of $j_{\gamma\gamma^{\prime}}^{\#}(\{N\})$ under the adjoint transformation $f_{\gamma\gamma^{\prime}}^{ch}\to-f_{\gamma\gamma^{\prime}}^{ch}$. Similarly, $j_{i}^{sp}$ is the anti-symmetric part of $j_{i}^{\vec{\mathcal{D}}\#}(\{N\})$ under the adjoint transformation $\vec{f}_{i}^{sp}\to-\vec{f}_{i}^{sp}$. We have dropped the #\# and $\{N\}$ index for brevity. The exact expressions for $j_{\gamma\gamma^{\prime}}^{ch}$ and $j_{i}^{sp}$ read:
@@
$j_{\gamma\gamma^{\prime}}^{ch}$ and $j_{i}^{sp}$ in eq. 22 are the dissipative currents generated by non-conservative driving forces $f_{\gamma\gamma^{\prime}}$ and $\vec{f}_{i}^{sp}$, compared to the total reactive/diffusive currents $j_{\gamma\gamma^{\prime}}^{\#}(\{N\})$ and $j_{i}^{\vec{\mathcal{D}}\#}(\{N\})$ in eq. 16 that include currents generated by conservative forces. The total bidirectional current, commonly called traffic, is obtained by replacing - with ++ in eq. 8, which also equivalently quantifies the scaled fluctuations resulting from the transitions. The form of eq. 22 reveals that the amplitude of $j_{\gamma\gamma^{\prime}}^{ch}$ and $j_{i}^{sp}$, and subsequently $\langle\dot{\sigma}^{ch}\rangle$ and $\langle\dot{\sigma}^{sp}\rangle$, is proportional to the traffic (scaled fluctuations) in the direction orthogonal to the external non-conservative driving force.
@@
In eq. 23, we have decomposed the EPR into its four linearly independent orthogonal contributions. First, $d_{t}\psi_{\varepsilon}$ quantifies the rate of change of the free energy attributed to the external driving work required to change the control parameters $\{\lambda\}$ of $\varepsilon$. Second, $d_{t}D^{KL}_{\varepsilon}(t)$ quantifies the EPR due to the relaxation towards the Boltzmann distribution $P_{\{N\}}^{\varepsilon}$, where $d_{t}\psi_{\varepsilon}$ and $d_{t}D^{KL}_{\varepsilon}(t)$ are physically interpreted as the boundary terms in the $\{\lambda\}$ and $P_{\{N\}}$ spaces, respectively. Third, $\langle\dot{\sigma}^{nr}\rangle$ quantifies the EPR due to the anti-symmetric forces (and vorticity currents generated by them) between the non-reciprocal particles. Fourth, $\langle\dot{\sigma}^{sp}\rangle+\langle\dot{\sigma}^{ch}\rangle$ quantifies the EPR due to the external non-conservative forces along the transitions, a thermodynamic cost supported by an external enzymatic or colloidal reservoir, which generates the dissipative transition currents [107].
One identifies the non-adiabatic EPR $\langle\dot{\sigma}_{\varepsilon}^{na}\rangle=-d_{t}D^{KL}_{\varepsilon}(t)$ and the remaining terms as the housekeeping EPR $\langle\dot{\sigma}_{\varepsilon}^{hs}\rangle$. The orthogonal decomposition is more fundamental because it does not require the existence of a steady state, as $\varepsilon$ is well-defined for any dynamic state. The lower bound on the total EP is obtained using $\langle\dot{\sigma}_{\varepsilon}^{na}\rangle$ [106, 103], and it reads $\langle\sigma_{\varepsilon}^{na}\rangle\geq D^{KL}_{\varepsilon}(0)-D^{KL}_{\varepsilon}(\tau)\geq D^{KL}(P(0)||P(\tau))$. Other notable consequences of the orthogonal decomposition have utilized different reference gauges [108, 109, 110, 111, 112, 113, 114, 115]. Choosing $P_{\{N\}}^{ss}$ as the reference distribution, the nonadiabatic-housekeeping decomposition of $\langle\dot{\sigma}\rangle$ is detailed in appendix A.
@@ 2.4 Coarse-graining
We implement the Doi-Peliti coarse-graining (DPCG) procedure to obtain a coarse-grained description [116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129]. The DPCG procedure incorporates the fluctuations (discreteness of the microscopic particle nature) owing to the second-quantization approach used. The technical details of the coarse-graining procedure and its physical implications are summarized in Ref. [87].
@@ 3 Fluctuating Macroscopic Description
The set of coarse-grained mesostates is denoted by $\{N_{i}^{\#}\}$. Here, $\{N_{i}^{\#}\}$ is defined over the discrete (lattice) space for the mesoscopic description. Physically, it corresponds to lumping the microstates (particles) with the same dynamic and thermodynamic properties into a single extensive coarse-grained mesostate. It achieves the coarse-graining step from (a) to (b) shown in fig. 1. We define $\Omega=N_{tot}/\mathcal{V}$, where $N_{tot}$ and $\mathcal{V}$ are the total number of particles in the whole lattice and the volume of the lattice (number of lattice sites), respectively. Thus, $\Omega$ quantifies the mean number of particles per lattice site. $N_{i}^{\#}$ is $O(\Omega)$ and its fluctuations due to microscopic transitions are $O(1)$ [130, 87].
To proceed further with the mesoscopic to macroscopic coarse-graining step, (b) to (c) depicted in fig. 1, we introduce the mesoscopic to macroscopic scaling $N_{i}^{\#}=\Omega\rho_{i}(\vec{x})$ between the macrostate $\rho_{i}(x)$ and the particle number mesostate $N_{i}^{\#}$. Therefore, $\rho_{i}$ and its fluctuations due to microscopic transitions are $O(1)$ and $O(1/\Omega)$ respectively [130, 87]. By construction, $\Omega\geq 1$, as $\Omega<1$ contradicts the mesoscopic to macroscopic coarse-graining. The space field $\vec{x}$ tracks the lattice index, which we drop in our notations and its meaning is presumed. Importantly, we use the macrostate convention consistent with the large-deviation theory [130, 87], where an intensive variable is defined and its fluctuations are suppressed with the large-deviation scaling parameter $\Omega$. Thus, the notion of an intensive variable is defined in the context of fluctuations. The macrostate $\rho_{i}$ does not necessarily coincide with the convention of the density used in hydrodynamics unless $\Omega=\mathcal{V}$ is chosen. This is fulfilled in the hydrodynamic limit, where the average number of particles per lattice site scales with the system volume. Hence, $N_{i}^{\#}=\mathcal{V}\rho_{i}$ holds. Importantly, this flexible definition of the macrostate $\rho_{i}$ allows us to define and study the systems on intermediate observation scales between the mesoscopic and hydrodynamic descriptions. $\Omega=1$ corresponds to the mesoscopic description that uses the particle number as the coarse-grained state. The thermodynamic limit of $\Omega\to\infty$ suppresses macrostate fluctuations and recovers the deterministic limit corresponding to the reaction-diffusion systems, fig. 1(d).